You have 18 milliliters of a 45% alcohol solution. How many milliliters of a 51% alcohol solution needs to be added to make a 49% alcohol solution?
Answer by stanbon(75887)

You have 18 milliliters of a 45% alcohol solution. How many milliliters of a 51% alcohol solution needs to be added to make a 49% alcohol solution?
------
Equation:
alcohol + alcohol = alcohol
0.45*18 + 0.51x = 0.49(18+x)
-------
Multiply thru by 100 to get:
45*18 + 51x = 49*18 + 49x
-----
2x = 4*18
x = 36 ml (amt of 51% solution to add)
=========================================
Cheers,
Stan H.
===========================